Main INDEX
Monthly INDEX
PREV
NEXT
User name xqian
Log entry time 11:26:37 on June16,2007
Entry number 205616
Followups:
keyword=target IOC reboot
The target IOC is rebooted due to lost of connection (a white alarm).
After the IOC reboot, the LOOP 3 temperatur alarm limit went wrong. Reset the higher limit to 120.0 and 130.0.
The FAN reading on the strip chart is 0. However, it is 60 in the main frame. It seems to be a strip chart reading problems. Target expert said it should be fine.
Now the target is all fine.